the combined ages of joel his father and grandfather are equal to 100. joel is 25 years younger then his dad,and 51 years younger then his grand father , how old is joel his father and his grandfather
Answer by ankor@dixie-net.com(22740)

the combined ages of joel his father and grandfather are equal to 100. joel is 25 years younger then his dad,and 51 years younger then his grand father , how old is joel his father and his grandfather
:
j + f + g = 100
:
" joel is 25 years younger then his dad,"
j = f - 25
Rewrite equation to
f = j + 25
:
"and 51 years younger then his grand father"
j = g - 51
rewrite equation to
g = j + 51
:
In the first equation replace f and g
j + (j+25) + (j+ 51) = 100
:
You should be able to do this now




, how old is joel his father and his grandfather